OpenStudy (anonymous):

In 1957 the price of a McDonald’s hamburger was 43¢. In 2002, the price of a hamburger was 88¢. Assuming the price increased linearly, what was the price of a McDonald’s hamburger in 1971?

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

ok we can do this

OpenStudy (anonymous):

idea is to find the equation of the line between two points \((1957,43)\) and \((2002,83)\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first we find the slope, which is \[\frac{83-43}{2002-1957}=\frac{40}{45}=\frac{8}{9}\]

OpenStudy (shane_b):

The slope will be 1...I think you used the wrong 2002 value :)

OpenStudy (anonymous):

caught me!

OpenStudy (anonymous):

\[\frac{88-43}{2002-1957}=\frac{45}{45}=1\]

OpenStudy (anonymous):

in other words the burger increases by a penny a year

OpenStudy (anonymous):

1971 is 14 years after 1957, so if it goes up a penny a year the price has increased by 14 cents from 43 cents

OpenStudy (anonymous):

ok